INFRARED SPECTRA OF TWO ISOMERS OF THE OCS-C_{2}H_{2} AND OCS-C_{2}D_{2}

\underline{MAHIN AFSHARI}, M. DEHGHANY, J. N. OLIAEE, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, Canada; A.R.W. MCKEL-LAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, ON K1A 0R6, Canada.

Spectra of the weakly-bound OCS-C_{2}H_{2} and OCS-C_{2}D_{2} complexes in the region of the OCS ν_{1} fundamental (∼2062 cm^{−1}) are observed in a pulsed supersonic slit jet expansion probed with a tunable diode laser. For each complex two bands are observed and assigned to the near parallel and the T-shape isomers. The ground state rotational and centrifugal distortion parameters were previousely determined from microwave studies.^{a} ^{b} Analysis of the infrared spectra gives accurate band origins as well as rotational and centrifugal distortion parameters for the upper states. All four bands show a red shift with respect to the monomer band origin, with the T-shape isomers having about 5.4 cm^{−1}larger shift than the corresponding near parallel isomers.

aS.A. Peebles and R.L. Kuczkowski, J. Phys. Chem. A 103, 3884 (1999).

bS.A. Peebles and R.L. Kuczkowski, Chem. Phys. Lett. 312, 357 (1999).